# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

### v3

@@ 4.2 Evaluation on SPAR-Bench
Tab. 4 shows that human performance far exceeds all models across all difficulty levels. The average human score is 67.27, with 55.31 on low-level tasks, 72.32 on middle-level tasks, and 76.22 on high-level tasks. Among baseline models (excluding our fine-tuned version), Qwen2.5-VL-72B performs best, but still falls far below human level, with an average score of 39.40 (35.35 on low-level, 23.05 on middle-level, and 48.44 on high-level tasks). This highlights that even the strongest models struggle with complex spatial reasoning, particularly as task difficulty increases.
Limitations of existing models.
The overall performance trend indicates that current models lack strong spatial perception and generalization ability for higher-order reasoning. The significant drop in scores from low-level to middle-level tasks suggests that models struggle to transition from basic spatial perception to more complex spatial relationships. For example, GPT-4o drops from 29.25 (low-level) to 24.93 (middle-level), and Qwen2.5-VL-72B drops from 35.35 to 23.05, highlighting difficulties in handling increasing spatial complexity. Additionally, performance in high-level tasks remains low, with most models scoring below 50, showing that multi-view and global spatial reasoning remain major challenges.
@@
Fine-tuning InternVL2.5-8B on our dataset significantly improves performance across all levels. The fine-tuned model scores 65.53 on low-level, 63.01 on middle-level, and 61.29 on high-level tasks, showing notable gains in spatial reasoning. For example, Depth-OC increases from 25.78 (InternVL2.5-8B) to 81.53, and View Change Inference improves from 6.32 to 37.14. These improvements suggest that high-quality spatial data enhances model performance, particularly in multi-view reasoning. However, the persistent gap between models and human performance indicates that data augmentation alone is insufficient, and further advances in multi-view fusion and spatial representation learning are needed.
@@
Methods EM@1 BLEU-4 CiDEr
@@
Table 5: Comparison of SQA3D and ScanQA performance across different LLM-based models.
@@ 4.3 Evaluation on 3D-specific tasks
In addition to the previously discussed spatial benchmarks, there are also tasks specifically defined in 3D scenes, which typically involve questions related to the entire scene.
After pretraining the VLM using our dataset, we perform supervised fine-tuning (SFT) on 3D tasks to further investigate the spatial capabilities of the VLM.
@@
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in Tab. 5, despite not incorporating 3D information into the model, it still achieves the highest score when compared to the 3D LLMs.
@@
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the Tab. 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well. These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
@@
ViT SPAR Ratio 2D 3D Low Medium High
Frozen 60% 68.68 89.67 66.87 64.28 55.43
Unfrozen 60% 73.00 91.10 67.01 44.04 51.86
Unfrozen 40% 74.13 89.09 66.97 42.21 48.19
Table 7: Performance comparison of different ViT configurations. Models are trained with varying SPAR ratios and evaluated on CV-Bench (2D, 3D) and SPAR-Bench.
@@ 4.4 Ablation study
We evaluate the impact of ViT training strategies and SPAR data proportions on InternVL2.5-4B, trained with 2M samples from SPAR (3D) and EMOVA (2D). Tab. 7 presents results on CV-Bench (2D, 3D) and SPAR-Bench (Low, Middle, High levels).
Effect of freezing vs. unfreezing ViT.
Comparing the first two rows with 60% SPAR, unfreezing ViT improves CV-Bench 2D (68.68 to 73.00) and 3D (89.67 to 91.10), but significantly degrades SPAR-Bench Middle (64.28 to 44.04) and High (55.43 to 51.86). This suggests that fine-tuning ViT benefits general vision tasks but weakens spatial reasoning, particularly in multi-view scenarios.
Effect of SPAR-7M ratio.
Comparing the second and third rows with Unfrozen ViT, reducing the SPAR ratio from 60% to 40% slightly improves CV-Bench 2D (73.00 to 74.13) but reduces CV-Bench 3D (91.10 to 89.09). SPAR-Bench performance further drops, particularly in middle level (44.04 to 42.21) and high level (51.86 to 48.19) tasks, indicating that higher 3D data proportions enhance multi-view spatial reasoning.
Findings.
Freezing ViT preserves spatial understanding, while unfreezing benefits 2D tasks but harms multi-view reasoning. Higher SPAR ratios improve 3D reasoning, while increasing 2D data slightly benefits 2D vision at the cost of 3D perception.
@@ 5 Conclusion
We introduce SPAR-7M, a large-scale dataset constructed from thousands of scenes, specifically designed to enhance spatial reasoning. To provide a rigorous evaluation framework, we propose SPAR-Bench, a benchmark that offers a more comprehensive assessment of spatial understanding, supporting both single-view and multi-view inputs. Experimental results demonstrate that pretraining on SPAR-7M alongside general datasets enables models to achieve SOTA performance on VSI-Bench and CV-Bench 3D. Furthermore, fine-tuning on 3D task-specific datasets yields competitive results, highlighting the effectiveness of our dataset in bridging the gap between 2D and 3D spatial reasoning.

### v4

@@ 4.2 Evaluation on SPAR-Bench
Table 4 shows that human performance consistently exceeds all models across difficulty levels, with a clear margin. While the strongest baseline (Qwen2.5-VL-72B) achieves competitive results, it still falls far short of human-level performance—particularly in tasks requiring spatial reasoning across views or precise geometric estimation.
While performance across levels is not directly comparable due to different metrics (MRA vs. accuracy), comparing to human results reveals a clear trend: models perform moderately on low-level tasks (e.g., 35.4% for Qwen2.5-VL-72B), but drop significantly on middle-level tasks (23.1%), indicating difficulty in structured spatial reasoning. Performance partially rebounds on high-level tasks (48.4%), likely aided by language priors or pattern-based inference. Tasks involving viewpoint changes, spatial imagination, or indirect metric reasoning remain the most challenging, underscoring current limits in spatial generalization and multi-view consistency.
@@
Fine-tuning InternVL2.5-8B on SPAR-mix significantly boosts performance across all difficulty levels, reaching 65.53 (low), 63.01 (mid), and 61.29 (high). The largest gains occur in depth estimation and multi-view reasoning—e.g., View Change Inference improves from 6.32 to 37.14—demonstrating the effectiveness of structured spatial data in aligning geometric cues across viewpoints. Notably, the fine-tuned model surpasses human performance in depth prediction, and even the strongest pre-finetuning model exceeds humans on distance estimation. This is understandable, as these metrics are MRA-based and favor numerically optimized models over human intuition.
These results confirm that our dataset introduces valuable inductive bias for spatial reasoning. However, gains remain uneven, and a clear gap persists compared to human-level performance—highlighting the need for advances in multi-view fusion, spatial representation learning, and compositional reasoning.
@@
Methods 3D EM@1 BLEU-4 CiDEr
@@
Table 5: Comparison of SQA3D and ScanQA performance. “3D” indicates whether the model is infused with 3D information.
@@ 4.3 Evaluation on 3D-specific tasks
In addition to the previously discussed spatial benchmarks, there are also tasks specifically defined in 3D scenes, which typically involve questions related to the entire scene. After pretraining the VLM using our dataset, we perform supervised fine-tuning (SFT) on 3D tasks to further investigate the spatial capabilities of the VLM.
@@
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in table 6, despite not incorporating 3D information into the model, it still achieves the competitive scores when compared to the 3D-enhanced LLMs.
@@
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the table 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well.
These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
@@
SPAR Ratio 2D 3D Low Medium High
40% 74.13 89.09 66.97 42.21 48.19
60% 73.00 91.10 67.01 44.04 51.86
Table 7: Performance comparison of different SPAR ratios.
@@ 4.4 Ablation study
We evaluate the impact of SPAR data proportions on InternVL2.5-4B, trained with 2M samples from SPAR (3D) and EMOVA (2D), using an unfrozen ViT. We adopt ViT fine-tuning as it improves overall performance and better adapts to the diverse spatial structures in SPAR-7M. Table 7 presents results on CV-Bench (2D, 3D) and SPAR-Bench (Low, Middle, High levels).
Effect of SPAR ratio.
With an unfrozen ViT, reducing the SPAR-7M ratio from 60% to 40% leads to a slight improvement on CV-Bench 2D (73.00 → 74.13), but results in performance drops on CV-Bench 3D (91.10 → 89.09) and across all levels of SPAR-Bench—particularly on the middle (44.04 → 42.21) and high (51.86 → 48.19) tasks. This indicates while 2D benchmarks may benefit from increased exposure to EMOVA-style data, an excessive reduction in the SPAR-7M proportion can cause the model to forget generalizable spatial priors. In contrast, complex spatial reasoning—especially under multi-view or relational conditions—relies more heavily on training signals tailored for spatial understanding. Maintaining or increasing the proportion of SPAR-7M is thus critical for reinforcing geometric alignment and ensuring spatial consistency across diverse viewpoints.
@@ 5 Conclusion
We introduce SPAR-7M, a large-scale dataset constructed from thousands of scenes, specifically designed to enhance spatial reasoning. To provide a rigorous evaluation framework, we propose SPAR-Bench, a benchmark that offers a more comprehensive assessment of spatial understanding, supporting both single-view and multi-view inputs. Experimental results show that pretraining on SPAR-7M alongside general datasets enables models to achieve state-of-the-art performance on VSI-Bench and CV-Bench 3D. Furthermore, fine-tuning on 3D task-specific datasets yields competitive results, highlighting the effectiveness of our dataset in bridging the gap between 2D and 3D spatial reasoning. While our tasks span diverse spatial challenges, they are primarily static QA-based and do not yet capture continuous spatial reasoning required in embodied or interactive settings.